OpenStudy (anonymous):

Which is a factor of 18x2+ 3x – 6 ?

OpenStudy (anonymous):

18x^2 + 3x - 6 = 3(6x^2 + x - 2) = 3(2x - 1)(3x + 2) I guess (3x - 2) would be the factor even though its incorrect. Either way, the above form is

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

MY MEDAL

OpenStudy (anonymous):

dont know